OpenStudy (anonymous):

please help 2yx=7y+5 with point (-9,2) in slope intercept form

OpenStudy (vijay):

your points are not satisfy with equation.

OpenStudy (anonymous):

i'm trying to help a friend 300 miles from me... but it's been forever since i've taken a math class

OpenStudy (anonymous):

y=5/2x-7+9/5

OpenStudy (anonymous):

how did you come to that

OpenStudy (vijay):

Actually this equation looks parabola ( not straight line )

OpenStudy (anonymous):

well using the equation y=mx+b for the slope intercept form you solve for b by plugging in the point (-9,2) to the equation y=5/2x-7 the -9 is your x and 2 is your y value.

OpenStudy (anonymous):

i am assuming you want it to pass through the point (-9,2) or it would not work since at x=-9 y does not equal 2 as vijay said it would not satisfy the equation

OpenStudy (anonymous):

right it needs to pass through that point

OpenStudy (vijay):

Felee, It will not go through the point, and this a parabolic equation. If you draw the graph , u will understand more about it.

OpenStudy (anonymous):

so if i were to simplify that equation it would be y=5/2x-8.8 ?????

OpenStudy (anonymous):

no the the -7 is part of the mx it needs to stay under the 5 its part of the slope. all you can do to simplify in these case is turn the 9/5 into a decimal Y=5/(2x-7)+1.8

OpenStudy (anonymous):

ok thank you i will pass this info along to my friend :)

OpenStudy (vijay):

nooooooooooooo, mr, cal2 student u r wrong u cannot make this eqution to y = mx+c format because , this is not straight line

OpenStudy (anonymous):

Thank you all for your help I'll pass it along